Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2021-08-05

Observers: Matt Nicholl, Evan Ridley, Ben Gompertz
Support: Sara Muñoz, María Delgado, Raúl Gonzalez, Lluís Galbany
TAT:  Jacob Teffs,  Michael Fulton,  Janet Chen


EFOSC Calibrations:
Afternoon: pessto_calib
(Morning: pessto_bias)

Weather:
Start: humidity ~ 15%, wind ~ 4m/s, seeing ~ 1.1"
Middle:
 humidity ~ 11%, wind ~ 8m/s, seeing ~ 1.4"
End: humidity ~ 11%, wind ~ 12m/s, seeing ~ 1.4"

Comments (times in UT):
Image analysis done after first standard
05:50 - LOEM failed due to network problem at La Silla, persevering over teams - target acquisition via TO's phone!!

 UTTarget 
 TypeSetup SlitExposure (s)Airmass
Seeing 
Notes
EFOSC2
        
23:27LTT7379
std
gr11-13-16
1-1.5
-
1.3

00:07AT2021uusclassificationgr131.09001.91.2
00:31 AT2021rlg classificationgr131.09001.30.8 
00:54 AT2021ubv classificationgr131.09001.41.1 
01:15 ZTF21abmemav classificationgr131.03001.21.1 
01:26 AT2021umz classificationgr131.09001.00.9 
01:45 SN2015Lfollowupgr11-161.02700 each1.61.1The seeing increased to 1.7 during the gr16 exposure. 
03:41 AT2021sqgclassificationgr131.56001.01.4 
04:00 Feige110 stdgr11-13-161-1.5-1.41.2 
04:14 AT2021uyl classificationgr131.01500   
04:43 SN2021uds followupgr11-161.01500 each1.41.0  
 6.01 SN2021uqw followupgr111.0 2 x 18001.3  Can't access images to measure seeing, DIMM suggests  ~1.5 
7.19
 SN2021tyw followupgr11-16 1.01500 each1.41.2
La Silla connection restored
8.23
 AT2021uzz classificationgr131.0
9001.01.4
 
8.45
 AT2021umf classificationgr13 1.015001.5
1.5 
9.22
 AT2021uuh classificationgr13 1.06001.11.4
 
9.38
 AT2021uvj classificationgr131.0
9001.2
1.4
 
 9.56 VMA2 stdgr11-13-161-1.5 -1.3
  
Comments